Consider a probability tree for selecting two puppies without replacement. The joint probability of two branches of the probability tree where the first selection is P(F)= 2/7 and the second selection P(F/F))= 3/10 is ______.

a. P(FF)=(5/17)

b. P(FF)= (2/7) (3/10)

c. P(FF)= (2/7/3/10)

d. P(FF)=22
Answer #1

Using conditional probability formula,  

P(F | F) = P(FF ) / P(F)

So,

P(FF) = P(F | F) * P(F)

= (2/7) * (3/10)